[figure: 9.7 cm]

POSSIBILITY OF HELICAL RKKY INTERACTION IN THE

MAGNETIC MULTILAYERS

Zygmunt Ba̧k^{a}

aInstitute of Physics, Pedagogical University of Czȩstochowa, al Armii Krajowej 13/15, 42-201 Czȩstochowa, Poland

Direct observation of the helical, nocollinear spin density waves /SDW/ in the Fe/Cr multilayers raised the question of noncollinear interlayer coupling in theses systems [1]. We assume that the SDW polarization within the spacer layer varies as:

m⃗_{±} = μ_{o} [x⃗ · cos (\frac{2π}{a} z) ± y⃗sin (\frac{2π}{a} z)]  (1)

where x⃗ and y⃗ are the unit cectors along the x, and y axis respectively, while a is the lattice constant. Taking into account the SDW pinning at the interfaces and coupling between correlated spin polarizability oscillations [2,]as well as the lock-in effects, we show that the resulting interlayer coupling can induce perpendicular ordering of the magnetizations of the adjacent layers.

[1] R.S.Fishman:J.Phys. (Cond. Matt.)13, (2001) R235-R269.

[2] Z. Ba̧k, W. Gruhn : J. Mag. Mag. Mater. 135 (1994) L7.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : Z. Ba̧k

Address for correspondence : Instytut Fizyki WSP, al. Armii Krajowej 13/15, 42-201 Czȩstochowa

Email address : z.bak@wsp.czest.pl